OpenStudy (anonymous):

if water covers 1,911 square miles of a state and this is only 5.9% of the area of the state. How many total square miles is the state?

OpenStudy (anonymous):

if 5.9% =1,911sm 1%=1911/5.9 100%(total sm)=1911/5.9 * 100